OpenStudy (insa):

fifty-four solid hemispheres each of diameter 2cm, are melted to form a single sphere. find the radius of the sphere.

OpenStudy (paki):

volume of one sphere will be V= 4/3 * pie* r^3

OpenStudy (sidsiddhartha):

54 solid hemisphere=27 spheres now find the volume of these 27 spheres

OpenStudy (paki):

so volume of 54 54 solid hemisphere will be... = 54(4/3)(pi)2^3 = 576 pie

OpenStudy (insa):

how volume of those hemispheres gets us the radius of the other sphere??

OpenStudy (paki):

hmmm , easy... just use another formula with little simplification....

OpenStudy (paki):

and then by doing so, you will need the radius... which is equal to the "vol of single sphere" = "vol of the 54 solid hemisphere"... so put the valuse and solve... (4/3)(pi)r^3 = 576(pi) (4/3)r^3 = 576 r^3 = 576(3/4) r^3 = 144(3) r^3 = 432 take cube rrot and solve... r = 7.56 cm

OpenStudy (paki):

@insa

OpenStudy (sidsiddhartha):

i think it should (4/3)(pi)r^3 = 36(pi) then r=3 u cant take 576pi because its volume of hemisphere and u can't equate it with (4/3)(pi)r^3

OpenStudy (insa):

okay now i get it thx again :)

OpenStudy (paki):

pleasure...